OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1?

jimthompson5910 (jim_thompson5910):

|2x – 3| – 8 = –1 |2x – 3| – 8 +8 = –1+8 |2x – 3| = 7 2x – 3 = 7 or 2x – 3 = -7

jimthompson5910 (jim_thompson5910):

Solve each equation to find your two answers

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

ok thank u

jimthompson5910 (jim_thompson5910):

yw